# Gravitational Wave Spectral Shapes as a probe of Long Lived Right-handed Neutrinos, Leptogenesis and Dark Matter: Global versus Local BLB-L Cosmic Strings

### v1

@@ 2 $^{3} 1001$$^{0}$ $10^{3}$ $10^{2}$
Table 2: The benchmark cases presented in Fig.7 for the RHN sourced matter domination.
@@ III.2 Gravitational Wave Tests of Long-Lived right-handed Neutrinos
By correlating the results for the detectability of the modified gravitational wave spectrum from cosmic strings, we identify the experimentally testable regions of the temperature parameter space. The testable regions by future experiments in the temperature plane are shown in Figure 8. Figure 9 shows the testable regions characterised by the end of matter domination and by its duration, parametrised by $N_{e}$.
@@
Using our best fit formulas for the onset and termination of matter domination in terms of the seesaw physics parameters $M$ and $\tilde{m}$, these detectable ranges can be trivially converted to detectable regions of the mass parameter spaces instead of temperature. This detectable parameter space is shown in Figure 10, while the overall procedure correlating gravitational-wave observables to the underlying seesaw parameters is illustrated in Fig. 11.
@@
Figure 10: Detectable regions of the seesaw parameter space inferred from the gravitational-wave background, where both the breaking and domination frequencies are determined using the best-fit relations (Eqs. 31, 34). The panels show global (left) and local (right) symmetry cases shown in the mass plane. These regions allow us to determine the mass of the right-handed neutrino and the effective neutrino mass related to the EMD caused by the right-handed neutrino.
@@
In Fig. 8, a monotonic suppression of the SGWB amplitude is observed with decreasing vev. This suppression progressively erodes the sensitivity of GW detectors across their maximum operational frequency bands. Beyond a critical threshold in amplitude, the signal effectively becomes observationally inaccessible. Consequently, the combined reach of the detector network fails to capture the complete spectral feature across a significant portion of the parameter space. This dynamic results in the emergence of isolated, island-like regions of detectability, which are confined to progressively lower vevs. The Fig. 10 clearly shows that the gravitational waves from global cosmic strings can probe right-handed neutrino masses across a range of nine orders of magnitude, spanning from approximately $M\sim 10\,\mathrm{GeV}$ up to $M\sim 10^{10}\,\mathrm{GeV}$, with sensitivity down to effective masses of order $\tilde{m}\sim 10^{-9}\,\mathrm{eV}$. Nonetheless, local strings extend this reach considerably, covering right-handed neutrino masses from $M\sim 0.1\,\mathrm{GeV}$ scale to $10^{9}\,\mathrm{GeV}$, while probing effective mass as small as $\tilde{m}\sim 10^{-10}\,\mathrm{eV}$. Symmetry breaking with larger vevs leads to a higher amplitude of the gravitational-wave background, while shifting the characteristic frequencies to much lower values, making them easier to detect. These results indicate that gravitational wave backgrounds from cosmic strings can explore a vast range of parameter space corresponding to such exotic matter domination by right-handed neutrinos [^2] If these strings are meta-stable meta1 ; meta2 ; meta5 ; meta7 , the PTA bound can be relaxed, allowing us to explore nearly the entire viable leptogenesis parameter space. Interestingly, in this case, an EMD is naturally motivated Datta:2024bqp to avoid the LIGO O3’s null result at higher frequencies.. In particular, the sensitivity to such small values of $\tilde{m}$ demonstrates that even extremely weakly coupled right-handed neutrinos, corresponding to lifetimes far beyond laboratory reach, leave an imprint that could be accessible to future GW detectors. Since the seesaw requires heavy Majorana neutrinos, mapping out this space is directly tied to testing the Majorana nature of neutrinos. Even laboratory searches, such as neutrino-less double beta decay [^3] From the experimental constraints of neutrinoless double $\beta$-decay from KATRIN, the direct neutrino mass measurement gives KATRIN:2021uub (see also Refs. Dolinski:2019nrj ; Gomez-Cadenas:2011oep ), $m_{\nu}\leq 0.8$ eV and future sensitivity can reach up to $m_{\nu}\leq 0.2$ eV., often regarded as hallmark laboratory tests of the seesaw mechanism, only constrain low‑energy effective parameters of the light neutrinos and cannot access the effective mass $\tilde{m}$ associated with heavy right‑handed states. When such states lie at scales beyond the reach of direct collider searches, gravitational waves may provide the only indirect and complementary probe of this crucial parameter of the type‑I seesaw. The impact of the presence of heavy RHN on the SM Higgs vacuum stability Ipek:2018sai has also been examined, but the associated bound is significantly higher than any sterile neutrino masses considered here and is therefore not constraining for our analysis. Another commonly discussed concern is Yukawa perturbativity. In our case, however, the effective neutrino masses are extremely small, ensuring that the Yukawa couplings remain well within the perturbative regime. For instance, even for $M=10^{12}$ GeV, the perturbativity bound corresponds to $\tilde{m}_{\rm max}\lesssim 400$ eV, far above the values considered here.
GWB observables $f_{\mathrm{brk}}$ $f_{\mathrm{dom}}$ $T_{\mathrm{dom}}$ $T_{\mathrm{end}}$ $M$ $\tilde{m}$ Eq. 48,50Eq. 31Eq. 4344M+M+ $T_{\mathrm{end}}+ Eq. 34$
@@
Figure 11: Flow chart from a gravitational-wave background (GWB) to seesaw parameters. A stochastic GWB can carry two key frequencies: $f_{\mathrm{dom}}$, marking the onset of matter domination, and $f_{\mathrm{brk}}$, corresponding to the return to radiation domination. From $f_{\mathrm{dom}}$, one infers the temperature of domination $T_{\mathrm{dom}}$ via Eqns. ((48),(50)), which determines the right-handed neutrino mass $M$ through Eq. 31. The break frequency $f_{\mathrm{brk}}$ gives the decay temperature $T_{\mathrm{end}}$ via Eqs. ((48),(50)). Finally, combining $M$ and $T_{\mathrm{end}}$ using Eq. 34 determines the effective neutrino mass $\tilde{m}$, providing a direct link between gravitational-wave observables and the seesaw parameters.
@@ IV Primordial Gravitational Wave Tests of Leptogenesis
The origin of the matter–antimatter asymmetry remains one of the central unresolved problems in cosmology and particle physics. The baryon asymmetry is measured independently from the cosmic microwave background Planck2018 and Big Bang Nucleosynthesis ParticleDataGroup:2020ssz
@@
Leptogenesis Fukugita:1986hr ; Luty:1992un offers a compelling and theoretically consistent framework that reproduces the observed baryon asymmetry: CP-violating decays of heavy right-handed neutrinos consistent with the Sakharov conditions Sakharov:1967dj generate a lepton asymmetry that electroweak sphalerons subsequently convert into a baryon asymmetry Luty:1992un ; Giudice_2004 ; Covi_1996 ; Buchm_ller_2005 . In the parameter regime considered here, where the Universe experiences a transient matter-dominated phase due to long-lived sterile neutrinos, the dynamics of leptogenesis are modified by both the altered expansion rate and subsequent entropy dilution. The goal of this section is to identify the regions in the $(M_{1},\tilde{m}_{1})$ parameter space that yield the observed baryon asymmetry and are potentially testable through cosmological signatures. The key parameter governing successful leptogenesis in both the hierarchical and near-resonant regimes is the right-handed neutrino mass $M_{1}$. As $M_{1}\propto v_{\rm B-L}$ originates from the spontaneous breaking of the $U(1)_{B-L}$ symmetry, the generation of the baryon asymmetry can ultimately be traced back to the dynamics of $U(1)_{B-L}$ breaking. The scale of this breaking also determines whether the gravitational-wave background can probe the period of matter domination through the detection of the characteristic frequencies $f_{\rm brk}$ and $f_{\rm dom}$.
@@ IV.1 The CP asymmetry
where $\Gamma$ represents the decay rate of the processes. The precise value of $\epsilon$ depends on the masses of the right-handed neutrinos and their Yukawa couplings Covi_1996 ; Buchm_ller_2005 ; Giudice_2004 ; Di_Bari_2012
@@
From this expression, we observe that $\epsilon$ becomes large when the masses of two right-handed neutrinos are nearly degenerate, $M_{i}\approx M_{j}$, leading to the phenomenon called resonant leptogenesis Pilaftsis_2004 . This scenario requires a refined treatment of the self-energy contribution, which dominates in this limit and develops a regulated enhancement. Various prescriptions for regulating the divergent behaviour have been proposed in the literature Klaric_2021 ; Pilaftsis_2004 ; Riotto_2007 ; Garbrecht_2014 ; Garny_2013 ; Anisimov_2006 . Owing to this ambiguity, in the present study, we refrain from analysing this regime. To this end, we shall impose the condition that the mass splitting is much greater than the decay rates of the right-handed neutrinos Moffat_2018 . Enhancements can occur through other mechanisms such as soft leptogenesis Boubekeur:2002jn ; D_Ambrosio_2003 ; however, they require extending the model with additional supersymmetric particles and soft-breaking terms, which goes beyond the minimal framework considered here.
@@ IV.2 Entropy Dilution and the Boltzmann Equations
Figure 13: Entropy dilution as a function of seesaw parameters. Left: Dependence on the effective neutrino mass. A smaller effective mass extends the period of matter domination, leading to the right-handed neutrino energy density dominating the energy budget to a larger extent. This gives a larger entropy dump when the right-handed neutrinos eventually decay. Right: dependence on both $M_{1}$ and $\tilde{m}$ showing entropy dilution is independent of right-handed neutrino mass.
@@
$\Delta=\left(1+3.726\times 10^{-6}\,\tilde{m}^{-0.67}\right)^{3/4}.$ (58)
The entropy dilution is heavily related to the duration of matter domination, since a prolonged phase enables right-handed neutrinos to overtake the energy content more substantially, which amplifies the entropy injection when they eventually decay.
To calculate the baryon asymmetry, we solve the Boltzmann equations for the lepton asymmetry as well as for the other variables. The asymmetry equation has two contributions: a source term, which generates the asymmetry, and a washout term, which accounts for inverse decays and $\Delta L=2$ scatterings Fong_2012 ; Ulysses ; Ulysses2 .
@@
$D(z)$ is the decay term in Eq. 30, the CP asymmetry parameter, $\epsilon$, was given in Eq. 55. $W(z)$ describes the washout. In the large $z$ limit, the equilibrium number density is exponentially suppressed Moffat_2018 ; Ulysses ; Ulysses2 ,
@@
and therefore, we have a vanishing washout term Moffat_2018 ; Ulysses ; Ulysses2
@@
In this regime, there is negligible erasure of the produced asymmetry. Moreover, with no efficient washout interactions, flavour effects Nardi:2006fx ; Abada_2006 ; Antusch_2006 ; Blanchet_2007 ; DeSimone:2006nrs ; Cirigliano_2010 ; Simone_2007 ; Racker_2012 are absent: the asymmetry generated in each lepton flavour evolves identically, so the single unflavoured Boltzmann equation is the appropriate equation to evolve the asymmetry. Temperature effects Giudice_2004 can also be omitted as leptogenesis in the regime $M\gg T$ where these effects are negligible. The final baryon asymmetry is then proportional to $Y_{B-L}$ through the sphaleron conversion factor and inversely proportional to the entropy dilution factor

### v2

@@ 2 $^{3} 1001$$^{0}$ $10^{3}$ $10^{2}$
Table 2: The benchmark cases presented in Fig.7 for the RHN sourced early matter domination.
@@ III.1.1 Gravitational Wave Detectors
  • Ground-based interferometers: These detectors, such as LIGO/VIRGO Abbott and others (2016b, a, 2017f, 2017a, 2017b, 2017c), aLIGO/aVIRGO Aasi and others (2015a); Acernese and others (2015); Abbott and others (2021), AION Badurina et al. (2021); Graham et al. (2016, 2017); Badurina and others (2020), Einstein Telescope (ET) Punturo and others (2010); Hild and others (2011), and Cosmic Explorer (CE) Abbott and others (2017e); Reitze and others (2019), use interferometric techniques on the Earth’s surface to detect gravitational waves.
  • Space-based interferometers: Space-based detectors like LISA Baker and others (2019), BBO Crowder and Cornish (2005); Corbin and Cornish (2006); Cutler and Holz (2009), DECIGO, U-DECIGO Seto et al. (2001); Yagi and Seto (2011), AEDGE El-Neaj and others (2020); Badurina et al. (2021), and $\mu$-ARES Sesana and others (2021) are designed to detect gravitational waves from space, offering different advantages over ground-based counterparts.
  • Recasts of star surveys: Monitoring of star surveys like GAIA/THEIA Garcia-Bellido et al. (2021) utilize astrometric data from stars can indirectly infer the presence of gravitational wave signals.
  • Pulsar timing arrays (PTA): PTA experiments like SKA Carilli and Rawlings (2004); Janssen and others (2015); Weltman and others (2020), EPTA Lentati and others (2015); Babak and others (2016), and NANOGRAV Arzoumanian and others (2018); Aggarwal and others (2019); Arzoumanian and others (2020) use precise timing periodicity measurements of pulsars to detect gravitational wave signatures.
@@ III.2 Gravitational Wave Tests of Long-Lived right-handed Neutrinos
By correlating the results for the detectability of the modified gravitational wave spectrum from cosmic strings, we identify the experimentally testable regions of the temperature parameter space. The testable regions by future experiments in the temperature plane are shown in Figure 8. Figure 9 shows the testable regions characterised by the end of early matter domination and by its duration, parametrised by $N_{e}$.
@@
Using our best fit formulas for the onset and termination of early matter domination in terms of the seesaw physics parameters $M$ and $\tilde{m}$, these detectable ranges can be trivially converted to detectable regions of the mass parameter spaces instead of temperature. This detectable parameter space is shown in Figure 10, while the overall procedure correlating gravitational-wave observables to the underlying seesaw parameters is illustrated in Fig. 11.
Figure 10: Detectable regions of the seesaw parameter space inferred from the gravitational-wave background, where both the breaking and early-matter domination frequencies are determined using the best-fit relations (Eqs. 31, 34). Detectability is defined by requiring that both characteristic frequencies are accessible within the combined sensitivity envelope of future GW detectors
@@
. The panels show global (left) and local (right) symmetry cases shown in the mass plane. These regions allow us to determine the mass of the right-handed neutrino and the effective neutrino mass related to the EMD caused by the right-handed neutrino.
@@
In Fig. 8, a monotonic suppression of the SGWB amplitude is observed with decreasing vev. This suppression progressively erodes the sensitivity of GW detectors across their maximum operational frequency bands. Beyond a critical threshold in amplitude, the signal effectively becomes observationally inaccessible. Consequently, the combined reach of the detector network fails to capture the complete spectral feature across a significant portion of the parameter space. This dynamic results in the emergence of isolated, island-like regions of detectability, which are confined to progressively lower vevs. The Fig. 10 clearly shows that the gravitational waves from global cosmic strings can probe right-handed neutrino masses across a range of nine orders of magnitude, spanning from approximately $M\sim 10\,\mathrm{GeV}$ up to $M\sim 10^{10}\,\mathrm{GeV}$, with sensitivity down to effective masses of order $\tilde{m}\sim 10^{-9}\,\mathrm{eV}$. This broad reach in $M$ arises from varying the Yukawa couplings that relate the right-handed neutrino masses to the underlying symmetry-breaking scale, $M=y_{N}v_{\rm B-L}$, while the gravitational-wave signal itself is primarily controlled by $v_{\rm B-L}$. Consequently, sizeable hierarchies between $v_{\rm B-L}$ and $M$ are both technically natural and phenomenologically relevant. Nonetheless, local strings extend this reach considerably, covering right-handed neutrino masses from $M\sim 0.1\,\mathrm{GeV}$ scale to $10^{9}\,\mathrm{GeV}$, while probing effective mass as small as $\tilde{m}\sim 10^{-10}\,\mathrm{eV}$. Symmetry breaking with larger vevs leads to a higher amplitude of the gravitational-wave background, while shifting the characteristic frequencies to much lower values, making them easier to detect. These results indicate that gravitational wave backgrounds from cosmic strings can explore a vast range of parameter space corresponding to such exotic early matter domination by right-handed neutrinos [^4] If these strings are meta-stable Vilenkin (1982); Preskill and Vilenkin (1993); Buchmuller et al. (2020, 2023), the PTA bound can be relaxed, allowing us to explore nearly the entire viable leptogenesis parameter space. Interestingly, in this case, an EMD is naturally motivated Datta and Samanta (2025a) to avoid the LIGO O3’s null result at higher frequencies.. In particular, the sensitivity to such small values of $\tilde{m}$ demonstrates that even extremely weakly coupled right-handed neutrinos, corresponding to lifetimes far beyond laboratory reach, leave an imprint that could be accessible to future GW detectors. Since the seesaw requires heavy Majorana neutrinos, mapping out this space is directly tied to testing the Majorana nature of neutrinos. Even laboratory searches, such as neutrino-less double beta decay [^5] From the experimental constraints of neutrinoless double $\beta$-decay from KATRIN, the direct neutrino mass measurement gives Aker and others (2022) (see also Refs. Dolinski et al. (2019); Gomez-Cadenas et al. (2012)), $m_{\nu}\leq 0.8$ eV and future sensitivity can reach up to $m_{\nu}\leq 0.2$ eV., often regarded as hallmark laboratory tests of the seesaw mechanism, only constrain low‑energy effective parameters of the light neutrinos and cannot access the effective mass $\tilde{m}$ associated with heavy right‑handed states. When such states lie at scales beyond the reach of direct collider searches, gravitational waves may provide the only indirect and complementary probe of this crucial parameter of the type‑I seesaw. The impact of the presence of heavy RHN on the SM Higgs vacuum stability Ipek et al. (2018) has also been examined, but the associated bound is significantly higher than any sterile neutrino masses considered here and is therefore not constraining for our analysis. Another commonly discussed concern is Yukawa perturbativity. In our case, however, the effective neutrino masses are extremely small, ensuring that the Yukawa couplings remain well within the perturbative regime. For instance, even for $M=10^{12}$ GeV, the perturbativity bound corresponds to $\tilde{m}_{\rm max}\lesssim 400$ eV, far above the values considered here.
GWB observables $f_{\mathrm{brk}}$ $f_{\mathrm{dom}}$ $T_{\mathrm{dom}}$ $T_{\mathrm{end}}$ $M$ $\tilde{m}$ Eq. 54,56Eq. 31Eq. 4950M+M+ $T_{\mathrm{end}}+ Eq. 34$
@@
Figure 11: Flow chart from a gravitational-wave background (GWB) to seesaw parameters. A stochastic GWB can carry two key frequencies: $f_{\mathrm{dom}}$, marking the onset of early matter domination, and $f_{\mathrm{brk}}$, corresponding to the return to radiation domination. From $f_{\mathrm{dom}}$, one infers the temperature of domination $T_{\mathrm{dom}}$ via Eqns. ((54),(56)), which determines the right-handed neutrino mass $M$ through Eq. 31. The break frequency $f_{\mathrm{brk}}$ gives the decay temperature $T_{\mathrm{end}}$ via Eqs. ((54),(56)). Finally, combining $M$ and $T_{\mathrm{end}}$ using Eq. 34 determines the effective neutrino mass $\tilde{m}$, providing a direct link between gravitational-wave observables and the seesaw parameters.
@@ IV Primordial Gravitational Wave Tests of Leptogenesis
The origin of the matter–antimatter asymmetry remains one of the central unresolved problems in cosmology and particle physics. The baryon asymmetry is measured independently from the cosmic microwave background Collaboration (2020) and Big Bang Nucleosynthesis Zyla and others (2020)
@@
Leptogenesis Fukugita and Yanagida (1986); Luty (1992) offers a compelling and theoretically consistent framework that reproduces the observed baryon asymmetry: CP-violating decays of heavy right-handed neutrinos consistent with the Sakharov conditions Sakharov (1967) generate a lepton asymmetry that electroweak sphalerons subsequently convert into a baryon asymmetry Luty (1992); Giudice et al. (2004); Covi et al. (1996); Buchmüller et al. (2005). In the parameter regime considered here, where the Universe experiences a transient matter-dominated phase due to long-lived sterile neutrinos, the dynamics of leptogenesis are modified by both the altered expansion rate and subsequent entropy dilution. The goal of this section is to identify the regions in the $(M_{1},\tilde{m}_{1})$ parameter space that yield the observed baryon asymmetry and are potentially testable through cosmological signatures. The key parameter governing successful leptogenesis in both the hierarchical and near-resonant regimes is the right-handed neutrino mass $M_{1}$. As $M_{1}\propto v_{\rm B-L}$ originates from the spontaneous breaking of the $U(1)_{B-L}$ symmetry, the generation of the baryon asymmetry can ultimately be traced back to the dynamics of $U(1)_{B-L}$ breaking. The scale of this breaking also determines whether the gravitational-wave background can probe the period of early matter domination through the detection of the characteristic frequencies $f_{\rm brk}$ and $f_{\rm dom}$.
@@ IV.1 The CP asymmetry
where $\Gamma$ represents the decay rate of the processes. The precise value of $\epsilon$ depends on the masses of the right-handed neutrinos and their Yukawa couplings Covi et al. (1996); Buchmüller et al. (2005); Giudice et al. (2004); Di Bari (2012)
@@
From this expression, we observe that $\epsilon$ becomes large when the masses of two right-handed neutrinos are nearly degenerate, $M_{i}\approx M_{j}$, leading to the phenomenon called resonant leptogenesis Pilaftsis and Underwood (2004). This scenario requires a refined treatment of the self-energy contribution, which dominates in this limit and develops a regulated enhancement. Various prescriptions for regulating the divergent behaviour have been proposed in the literature Klaric et al. (2021); Pilaftsis and Underwood (2004); Simone and Riotto (2007a); Garbrecht et al. (2014); Garny et al. (2013); Anisimov et al. (2006). Owing to this ambiguity, in the present study, we refrain from analysing this regime. To this end, we shall impose the condition that the mass splitting is much greater than the decay rates of the right-handed neutrinos Moffat et al. (2018). Enhancements can occur through other mechanisms such as soft leptogenesis Boubekeur (2002); D’Ambrosio et al. (2003); however, they require extending the model with additional supersymmetric particles and soft-breaking terms, which goes beyond the minimal framework considered here.
@@ IV.2 Entropy Dilution and the Boltzmann Equations
Figure 13: Entropy dilution as a function of seesaw parameters. Left: Dependence on the effective neutrino mass. A smaller effective mass extends the period of early matter domination, leading to the right-handed neutrino energy density dominating the energy budget to a larger extent. This gives a larger entropy dump when the right-handed neutrinos eventually decay. Right: dependence on both $M_{1}$ and $\tilde{m}$ showing entropy dilution is independent of right-handed neutrino mass.
@@
$\Delta=\left(1+3.726\times 10^{-6}\,\left(\frac{\tilde{m}}{\rm eV}\right)^{-0.67}\right)^{3/4}.$ (64)
The entropy dilution is heavily related to the duration of early matter domination, since a prolonged phase enables right-handed neutrinos to overtake the energy content more substantially, which amplifies the entropy injection when they eventually decay.
To calculate the baryon asymmetry, we solve the Boltzmann equations for the lepton asymmetry as well as for the other variables. The asymmetry equation has two contributions: a source term, which generates the asymmetry, and a washout term, which accounts for inverse decays and $\Delta L=2$ scatterings Fong et al. (2012); Granelli et al. (2021, 2023).
@@
$D(z)$ is the decay term in Eq. 30, the CP asymmetry parameter, $\epsilon$, was given in Eq. 61. $W(z)$ describes the washout. In the large $z$ limit, the equilibrium number density is exponentially suppressed Moffat et al. (2018); Granelli et al. (2021, 2023),
@@
and therefore, we have a vanishing washout term Moffat et al. (2018); Granelli et al. (2021, 2023)
@@
In this regime, there is negligible erasure of the produced asymmetry. Moreover, with no efficient washout interactions, flavour effects Nardi et al. (2006); Abada et al. (2006); Antusch et al. (2006); Blanchet and Bari (2007); De Simone and Riotto (2007); Cirigliano et al. (2010); Simone and Riotto (2007b); Racker et al. (2012); Roshan and Show (2025) are absent: the asymmetry generated in each lepton flavour evolves identically, so the single unflavoured Boltzmann equation is the appropriate equation to evolve the asymmetry. Temperature effects Giudice et al. (2004) can also be omitted as leptogenesis in the regime $M\gg T$ where these effects are negligible. The final baryon asymmetry is then proportional to $Y_{B-L}$ through the sphaleron conversion factor and inversely proportional to the entropy dilution factor